OpenStudy (anonymous):

approximate the real root of 2x=1/e^x

OpenStudy (anonymous):

there are a number of algorithms for this...one simple one is to write it as x = 1/(2e^x) and the algorithm is x_n = 1/2 * 1/e^(x_n-1) so guess x_0 = 1: x_1 = 1/2 so x_2 = 1/2 * 1/e^(-1/2) = 1/2 * e^(1/2) so x_3 = 1/2 * 1/e^(1/2*e^(1/2) - 1) so after the third iteration we get an approximation of x = 0.5960

OpenStudy (anonymous):

oh wait...oops, lemme try again: x_1 = 1/2e x_2 = 1/2 * 1/e^(1/2e) = 0.4160 x_3 = 1/2 * 1/e^(0.4160) = 0.3298 x_4 = 1/2 * 1/e^(0.3298) = 0.3595 and so on

OpenStudy (anonymous):

It's a multiple choice question and these are the possible answers .352, .335, .352, .355, or 3.521

OpenStudy (anonymous):

x_5 = 1/2 * 1/e^(0.3595) = 0.3490 x_6 = 1/2 * 1/e^(0.3490) = 0.3527 x_7 = 1/2 * 1/e^(0.3527) = 0.3514

OpenStudy (anonymous):

x_9 = 0.3517

OpenStudy (anonymous):

so after 9 iterations, the approximation seems to be roughly 0.3517, which is roughly 0.352

OpenStudy (anonymous):

you can also use the newton-rhapsodian algorithm to approximate this the exact solution is given by W(1/2), where W is the lambert W function

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

did you need to use an algorithm or could you just enter this into a calculator? LOL

OpenStudy (anonymous):

I think it could've been entered into a calculator, but my teacher wants us to show all our work

OpenStudy (anonymous):

using newton-rhapson: x_n+1 = x_n - f(x_n)/f'(x_n) so x_n+1 = x_n - (2x_ne^x_n - 1)/2e^x_n(x_n+1) if we guess x_0 = 1: x_1 = 0.5920 x_2 = 0.3939 x_3 = 0.3532 x_4 = 0.3517 that's a lot faster

OpenStudy (anonymous):

Yeah, that is a lot faster, thanks for the help